# Nonasymptotic Analysis of Accelerated Methods With Inexact Oracle Under Absolute Error Bound

### v1

@@ 3.1 PEP approach for the convergence analysis of iGOGM
Surprisingly, this solution matches the numerical solution when $b_{i}\equiv b$. In Figure 3, we follow the same settings as for Figure 2 with $\mathbf{v}$ equals to the values defined in Equation 10. The error between the optimal solution and the analytical $\hat{\mathbf{u}}$ given in (12) is under $10^{-3}$, validating our theoretical result.
@@
Finally, in Lemma 3.2, we provide an analytical solution for $\mathbf{u}$ that together with the analytical solutions for $\mathbf{v}$ and $\mathbf{v}_{*}$ (from the problem with exact oracle) provide a feasible solution to (D).
From the above discussion, we can summarize the results in the following lemma. The lemma provides a feasible solution to problem (D) with iGOGM algorithm.
@@
For Algorithm 5, the following solution is a feasible solution to the problem (D). $\forall i=0,\cdots,K-1$
@@
$u_{i}$ $=\frac{A_{i}(1+2\alpha_{i+1})(A_{i}+2\alpha_{i}\alpha_{i+1})}{4LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}+\sum_{k=i+1}^{K-1}\frac{A_{k}(1+2\alpha_{k+1})\alpha_{i}\alpha_{k+1}}{2LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}.$
3.2 PEP-inspired convergence analysis of iGOGM
As explained by [30], the feasible solution of PEP provides the proof of the convergence bound for the first-order algorithms. In the following theorem and the corresponding proof, we will show a PEP-inspired approach that can be verified without any knowledge of PEP or even SDP.
Theorem 3.3.
The sequence generated by Algorithm 5 satisfies
Proof.
Below are the inequalities used in PEP primal constraints and their corresponding dual variables:
@@ 3.2 PEP-inspired convergence analysis of iGOGM
Multiply inequalities with their corresponding dual variable and sum up, we have the inequality
@@
$A_{K}(f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2})$
$\leq$ $-\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
$-\frac{1}{2L}\sum_{k=0}^{K-1}A_{k}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{A_{K}}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}(A_{k+1}\theta_{k+1,i}-A_{k}\theta_{k,i})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k+1}\theta_{k+1,k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mua\mkern-1.5mu)}}{{=}}$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
@@
$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mui\mkern-1.5mu)}}{{=}}$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau}-\frac{A_{k}}{L}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\alpha_{k}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mub\mkern-1.5mu)}}{{=}}$ $-\frac{4}{L}\left\|\frac{L}{4}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{4}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle,$
where (a) replaces $\theta_{k,i}$ with (11), and (b) uses $\tau=L/4$ and the fact that $A_{0}=\alpha_{0}=1$. If $\mathbf{e}_{i}=\mathbf{0}$, then the above inequality gives the convergence rate of iGOGM with exact oracle.
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}^{*}\right\|}{4A_{K}}$
Next, we will bound the error terms when $\mathbf{e}_{i}\neq\mathbf{0}$. For any $k=0,\cdots,K-1$,
@@
$-\frac{A_{k+1}-\alpha_{k+1}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{2}{L}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5muii\mkern-1.5mu)}}{{=}}$ $\frac{-1}{L(A_{k+1}-\alpha_{k+1}^{2})}\left\|(A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+\frac{1}{2}A_{k}\mathbf{e}_{k}\right\|^{2}$
$+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}^{2}\alpha_{i}^{2}\left\|\mathbf{e}_{i}\right\|^{2}+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k}\sum_{j=0}^{i-1}2\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle$
$+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}A_{k}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle+\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}.$
Summing over $k$, and changing the summation indices, we have
@@
$\sum_{k=0}^{K-1}\left(-\frac{A_{k+1}-\alpha_{k+1}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{2}{L}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle\right)$
$\leq$ $\sum_{k=0}^{K-1}\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{L(A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
@@
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{L(A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$

### v2

@@ 4 Generalized OGM with inexact gradient oracle
Surprisingly, this solution is close to the numerical solution when $b_{i}\equiv b$. In Figure 7, we follow the same settings as for Figure 6 with $\mathbf{v}$ equals to the values defined in (15). The error between the optimal solution and the analytical $\hat{\mathbf{u}}$ given in (19) is under $10^{-3}$, validating our theoretical result.
@@
We now summarize the above results in the following lemma. Specifically, Lemma 4.1 establishes a feasible solution to problem (D) for the iGOGM algorithm. This is achieved by providing an analytical solution for $\mathbf{u}$ which, when combined with the analytical solutions for $\mathbf{v}$ and $\mathbf{v}_{*}$ from the problem with the exact oracle, satisfies the problem’s constraints.
@@
Under Assumption 2.1, for the iGOGM (Algorithm 2), the following solution is a feasible solution to the problem (D).
@@
$u_{i}$ $=\frac{A_{i}(1+2\alpha_{i+1})(A_{i}+2\alpha_{i}\alpha_{i+1})}{4LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}+\sum_{k=i+1}^{K-1}\frac{A_{k}(1+2\alpha_{k+1})\alpha_{i}\alpha_{k+1}}{2LA_{K}(A_{k+1}-\alpha_{k+1}^{2})},\forall i=0,\cdots,K-1.$
4.1 PEP-inspired proof of the convergence analysis of iGOGM
In this section, we provide the detailed proof of Theorem 2.2. As explained in [25], a feasible solution to the PEP’s dual problem (SDP-PEP), or its relaxed version (D), provides a direct proof for the algorithm’s convergence bound. We note that even though we use the PEP technique to find an upper bound on the convergence bound, the proof can be understood without any prior knowledge of the PEP or SDP’s duality theory.
Following this idea, we will establish our proof for iGOGM based on the feasible solution we derive in Lemma 4.1. Recall the properties of Lagrangian duality, any such feasible solution of (D) yields the following upper bound on the optimization error
$f(\mathbf{x}_{K})-f_{*}\leq$ $\mathcal{L}(\mathbf{G},\mathbf{F},\mathbf{v},\mathbf{v}_{*},\mathbf{u},\tau)$
$=$ $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}+\mathbf{F}\left(\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})\right)$
$-\operatorname{Tr}\left(\mathbf{G}\left(\sum_{i=0}^{K-1}v_{i,i+1}\mathbf{A}^{i,i+1}+\sum_{i=0}^{K}v_{*,i}\mathbf{A}^{*,i}+\tau\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top}+\sum_{i=0}^{K-1}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}+\frac{1}{2L}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}\right)\right)$
$\leq$ $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}.$
Based on the above inequality, our proof is constructed by forming a linear combination of the inequalities derived from the function’s convexity and $L$-smoothness properties. The dual variables $\mathbf{v}$ and $\mathbf{u}$ act as the coefficients for this combination. The first step is to multiply the following inequalities by their corresponding dual variables
@@ 4.1 PEP-inspired proof of the convergence analysis of iGOGM
[…]
@@
Summing over $k$ and subtracting $\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ on both sides and rearranging terms we get
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$\leq$ $\tau(R^{2}-\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2})+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle$
$-\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle-\frac{1}{2L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}$
$-\frac{1}{2L}\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}.$
Replacing $\mathbf{x}_{k}$ with the update rule of iGOGM defined by (10) and (16), the above inequality becomes
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
@@
$\leq$ $\tau(R^{2}-\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2})+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle$
@@
$+\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}.$
To complete the proof, we must find $\tau$ and $\{u_{k}\}$ that ensures the right-hand-side of the above inequality is bounded by $\tau R^{2}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}$. This is equivalent to finding a feasible solution for the dual problem (D), since the bound is guaranteed if the positive semidefinite constraint (14) holds. Our procedure of finding $\mathbf{u}$ is exactly the step of canceling the remaining terms in the right-hand-side. We will provide a detailed explanation below.
@@
First, we do the Shur complement step corresponding to $\mathbf{M}=\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}+\mathbf{R}$, which is also the step to bound $\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle$.
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$\leq$ $\tau(R^{2}-\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2})+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\underbrace{\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}}_{\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}}$
@@
$+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau A_{K}^{2}}-\frac{A_{k}}{LA_{K}}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\frac{\alpha_{k}\alpha_{i}}{A_{K}^{2}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle$
@@
$+\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\frac{4A_{K}}{L}\left\|\frac{L}{4A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle,$
where in the last equality we replace $\tau=\frac{L}{4A_{K}}$ and $A_{0}=\alpha_{0}=1$. In the exact case, $b_{k}\equiv 0$ and $\mathbf{e}_{k}\equiv\mathbf{0}$, the inner-product terms in the above inequality are all zero, and all norm terms have negative coefficients. The convergence result of the exact GOGM can be recovered as
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}^{*}\right\|^{2}}{4A_{K}}.$
In the inexact case, we need to continue to find a value of $\{u_{k}\}$ that cancels the non-zero inner-product terms. To achieve this goal, we will follow the step of the Shur complement applied to matrix $\mathbb{R}^{\backslash[1,K+2]}$ (17), i.e., $\mathbf{R}^{\backslash[1,K+2]}/\mathbf{D}=\mathbf{U}-\mathbf{B}^{\backslash[1]}\mathbf{D}^{-1}\mathbf{B}^{\backslash[1]^{\top}}$:
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}^{2}\alpha_{i}^{2}\left\|\mathbf{e}_{i}\right\|^{2}$
$-\underbrace{\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\|(A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+\frac{1}{2}A_{k}\mathbf{e}_{k}\right\|^{2}}_{\mathbf{B}^{\backslash[1]}\mathbf{D}^{-1}\mathbf{B}^{\backslash[1]^{\top}}}$
$+\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k}\sum_{j=0}^{i-1}2\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle$
$+\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}A_{k}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle+\sum_{k=0}^{K-1}\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K-1}\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$